How much of a 30% mixture of fertilizer to be mixed w 70% fertilizer to get 40gallons of 60% fertilizer...how much 30% and how much 70%?
Answer by richwmiller(17219)

We need a total of 40 gal at 60%
We want to know how many gal at 30% to mix with 70%
a+b=40,
0.3*a+0.7*b=0.6*40
a=40-b
0.3*(40-b)+0.7*b=24
12-0.3b+0.7*b=24
0.4*b=12
b=30
a=40-b
a=10 gal at 30%
b=30 gal at 70%
check
0.3*10+0.7*30=0.6*40
3+21=24
24=24
ok
codemx1